OpenStudy (anonymous):

cscx cotx (1-cos^2x)= ____x Please explain how to solve this?

OpenStudy (debbieg):

Do you know an identity that allows you to substitute something in for (1-cos^2x)?

OpenStudy (debbieg):

\(\sin^2(x)+\cos ^2(x)=1\) might be helpful. :)

OpenStudy (anonymous):

I couldn't find an identity that matched that particular expression. How is the Pythagorean identity obtained from that?

OpenStudy (anonymous):

Wait... is the answer sinx? 1-cos^2x equal sin^2x, soooo.... eh?

OpenStudy (debbieg):

It isn't obtained from it. It's useful in simplifying it. Rearrange it and you'll get an expression for \(1-cos ^2(x)\)

OpenStudy (debbieg):

eh. :)

OpenStudy (debbieg):

Yes, after you make that substitution and convert the other factors to sines and cosines, everything just falls right into place, doesn't it? :)

OpenStudy (anonymous):

Hallelujah! I'm finally getting somewhere! Thank you so much. :)

OpenStudy (debbieg):

You're very welcome, happy to help. :)

OpenStudy (debbieg):

\[\Large \begin{matrix}\Uparrow ~~~~~~~~~~~~~~~~~~~~~~~~ \\ \heartsuit \text{'s trig identities}\end{matrix}\] Hehe.... don't mind me, just playing with LaTex. :)